OpenStudy (anonymous):

integral of (1+sinx)/(1-sinx)? seems like I should separate the integral but that doesn't seem to help much.

OpenStudy (anonymous):

multiply and divide with 1+sin(x) \[\Large \int\limits{}^{} \frac{1+\sin(x)}{1-\sin(x)}*\frac{1+\sin(x)}{1+\sin(x)}\] \[\int\limits \frac{(1+\sin(x))^2}{1-\sin^2(x)}dx=\int\limits \frac{1+\sin^2(x)+2\sin(x)}{\cos^2(x)}dx\]

OpenStudy (anonymous):

ok great thanks

OpenStudy (anonymous):

can you complete this or need further help with this ?

OpenStudy (anonymous):

Im fine with the other terms but Im having trouble with sin^2/cos^2?

OpenStudy (anonymous):

I mean just taking the integral of it.

OpenStudy (anonymous):

ok we can write this by distributing the denominator \[\int\limits (\frac{1}{\cos^2(x)}+\frac{\sin^2(x)}{\cos^2(x)}+\frac{2\sin(x)}{\cos^2(x)})dx\] which becomes \[\int\limits (\sec^2(x)+\tan^2(x)+2\tan(x)\sec(x))dx\] now seperate the integrals and solve them . you may know . \[\int\limits \sec^2(x)dx=\tan(x)\]

OpenStudy (anonymous):

yep ok got it thanks again.

OpenStudy (anonymous):

you welcome :)